Circular No. 8073

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003as IN MCG +08-10-7 Further to IAUC 8072, T. Boles reports the discovery of another ap-

parent supernova (mag 16.8) on an unfiltered CCD image taken on Feb. 13.909 and 14.765 UT. The new object is located at α = 5^{h}28^{m}45^{s}.81, δ = +49^{o}52′59′′.1, which is ≈ 10′′.1 east and 1′′.7 south of the center of MCG +08-10-7. Nothing was visible at this position on Boles’ images taken on Jan. 5 and 26 (limiting mag 19.0) and on Palomar Sky Survey red (1998 Oct. 8) or blue (1994 Dec. 3) plates.

SUPERNOVA 2003L IN NGC 3506 S. Kulkarni and D. W. Fox, California Institute of Technology, report

that a 30000-s exposure with the Chandra ACIS-S detector beginning at Feb. 10.154 UT detected 41 ± 6 counts from SN 2003L: “The best-fit power spectrum yields photon index 1.4 ± 0.4 for an assumed column density of 2.5 × 10^{20} cm^{−2} (2–10-keV flux of 6 × 10^{−15} erg cm^{−2} s^{−1}); equivalently, a blackbody of temperature 0.5 ± 0.1 keV (2–10-keV flux of 2 × 10^{−15} erg cm^{−2} s^{−1}) is also acceptable. The x-ray-to-radio luminosity ratio of SN 2003L is roughly half that observed for SN 1998bw at a similar epoch.”

COMETS C/2000 G3; C/2002 V3, V4, V5, X6, Y2 (SOHO) Further to IAUC 8065, D. Hammer reports measurements for six addi-

tional comets found on SOHO C2 website images by R. Kracht (C/2000 G3, C/2002 V4, C/2002 X6), D. Evans (C/2002 V3), and R. Matson (C/2002 V5, 2002 Y2). C/2000 G3, C/2002 V3, and C/2002 V5 were also visible in C3-coronagraph images. The reduced astrometry and orbital elements by B. G. Marsden appear on the MPECs cited below. C/2002 V4, C/2002 X6, and C/2002 Y2 are members of the Meyer group; C/2002 V5 is a member of the Marsden group; the others are Kreutz sungrazers.

Comet  2000 UT  α_{2000}  δ_{2000}  MPEC
C/2000 G3 Apr. 10.404  1^{h}27.^{m}1  + 5^{o}57′  2003-B42
Comet  2002 UT  α_{2000}  δ_{2000}  MPEC
C/2002 V3 Nov. 9.179  14^{h}50.^{m}3  −18^{o}28′  2003-B42
C/2002 V4  9.579  14 55.2  −15 36  2003-B42
C/2002 V5  11.854  15 05.5  −18 54  2003-C02
C/2002 X6 Dec. 2.314  16 32.7  −20 50  2003-C02
C/2002 Y2  19.579  17 51.0  −22 27  2003-C02
2003 February 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green